OpenStudy (anonymous):

a ball is thrown vertically upwards from a height of 6 ft with an initial velocity of 40 ft per second. how high will the ball go?

OpenStudy (anonymous):

well if we stay in ft and since we do not know a time (yet we can find this out without a time) we use the function \[v'^{2}=v^{2}-2gy\] We want to know y. We know g is 32ft/s^2, initial velocity (v) is 40 ft/s and at the top of the projection (max height) our final velocity (v') is zero correct?

OpenStudy (anonymous):

correct alll it wants is the max hieght

OpenStudy (zzr0ck3r):

is this for a calculus class on integration?

OpenStudy (anonymous):

yea its a calculus question

OpenStudy (zzr0ck3r):

if so, then we have \(\int g \ dt\) where \(g\) is a constant for gravity, and we also know \(v(0)=40\) and \(d(0)=6\)

OpenStudy (anonymous):

correct so insert the values \[0=200ft^{2}/s^{2}-2(32ft/s^{2})y\] \[-200ft^{2}/s^{2}=(-64ft/s^{2})y\] \[y=3.125ft\]

OpenStudy (anonymous):

and yea we need to include the initial height of 6ft so 9.125ft

OpenStudy (anonymous):

well i have possible answers and it doesn't have 3.125 ft

OpenStudy (anonymous):

here ill list the answers a. 81 ft, 24.7 ft c.31 ft d. 69 ft e. 84 ft

OpenStudy (zzr0ck3r):

\(v(t) = \int g \ dt= gt+c\\ v(0) = 40 = c\\v(t)=gt+40\\p(t)=\int v(t) dt=\frac{1}{2}gt^2+40t+c_0\\p(0)=6=c \\p(t)=\frac{1}{2}gt+40t+6\) The vertex \(t\) coordinate of the vertex is given by \(t=\frac{-40}{\frac{g}{2}}\) Plug in g for gravity (it should be negative) and then plug the result into \(p(t)\)

OpenStudy (zzr0ck3r):

What are you using for gravity?

OpenStudy (anonymous):

well the questions doesn't really give out the gravity

OpenStudy (zzr0ck3r):

That should say \(t=\frac{-40}{2(g\frac{1}{2})}=\frac{-40}{g}\) we will assume g = is -32 (negative because its working against us) \(p(\frac{-40}{-32})=p(\frac{5}{4})=31ft/s\)

OpenStudy (anonymous):

oh k, thanks zzr0ck3r for the help!

OpenStudy (zzr0ck3r):

31 ft not ft/sec...